Togo's Sandwich Shop had the following long-term asset balances as of January 1, 2021:

Accumulated
Cost Depreciation Book Value
Land $ 79,000 $ 79,000
Building 554,000 $ (199,440 ) 354,560
Equipment 143,600 (28,800 ) 114,800
Patent 110,000 (44,000 ) 66,000

Togo's purchased all the assets at the beginning of 2019 (3 years ago). The building is depreciated over a 10-year service life using the double-declining-balance method and estimating no residual value. The equipment is depreciated over a 9-year useful life using the straight-line method with an estimated residual value of $14,000. The patent is estimated to have a five-year service life with no residual value and is amortized using the straight-line method. Depreciation and amortization have been recorded for 2019 and 2020.

1. For the year ended December 31, 2021, record depreciation expense for buildings and equipment. Land is not depreciated. (If no entry is required for a transaction/event, select "No Journal Entry Required" in the first account field.)

------Record the depreciation on the building.

2. For the year ended December 31, 2021, record amortization expense for the patent. (If no entry is required for a transaction/event, select "No Journal Entry Required" in the first account field.)

------ Record the amortization on the patent.

3. Calculate the book value for each of the four long-term assets at December 31, 2021.
Answer #1

Answer 1 :

Date General Journal Debit Credit
December 31, 2021 Depreciation Expense-Building 70,912
Accumulated Depreciation-Building 70,912
(To record depreciation on buildings)
December 31, 2021 Depreciation Expense-Equipment 14,400
Accumulated Depreciation-Equipment 14,400
(To record depreciation on equipment)

Explanation:

Depreciation expense for Buildings :

SLM rate = 1 / 10 years = 10 %.

Depreciation as per double-declining-balance method = SLM rate * 2 * Book value at the beginning of the year

= 10 % * 2 * $354,560 = $70,912

Depreciation expense for Equipment :

Depreciation as per straight-line method = (Cost - Residual value) / Useful life

= ($143,600 - $14,000) / 9 years = $14,400.

Answer 2 :

Date General Journal Debit Credit
December 31, 2021 Patent Amortization 22,000
Patent 22,000
(To record amortization on patent)

Explanation:

Amortization expense for the patent = Cost / Useful life

= $110,000 / 5 years = $22,000.

Answer 3:

Book value for each of the four long-term assets at December 31, 2021:

Assets Book Value
Land $79,000
Building ($354,560 - $70,912) $283,648
Equipment ($114,800 - $14,400) $100,400
Patent ($66,000 - $22,000) $44,000